OpenStudy (anonymous):

w/4 – 4 = 3 –4 28 3 11 d/3+ 10 = 7 51 20 0 –9

OpenStudy (anonymous):

do you know?

OpenStudy (anonymous):

the first one is 3

OpenStudy (anonymous):

thx!

OpenStudy (anonymous):

do you know the second one?

OpenStudy (anonymous):

the second one is -9

OpenStudy (anonymous):

your a lifesaver thank you!

OpenStudy (anonymous):

lol ur welcome

OpenStudy (anonymous):

:)

OpenStudy (whpalmer4):

@thatgirlisheres Instead of just getting the answers, wouldn't you like to see *how* to get them yourself? \[\frac{w}{4}-3 = 3\]Multiply both sides by 4 to eliminate the fraction\[4*\frac{w}{4}-4*3=4*3\]\[w-12=12\]Add 12 to both sides to get w alone on the left side\]\[w-12+12=12+12\]\[w=24\] \[\frac{d}{3}+10=7\]Multiply both sides by 3 to eliminate the fraction\[3*\frac{d}{3}+10*3=7*3\]\[d+30=21\]Subtract 30 from both sides to get d alone on the left\[d+30-30=21-30\]\[d=-9\] See, not really so hard. You can do them yourself!

OpenStudy (anonymous):

wow thx i would like to see how but um no one explained it so yaaa

OpenStudy (whpalmer4):

Okay, I thought I did at least demonstrate how to do it. Do you think you understand now?

OpenStudy (anonymous):

ya thx